OpenStudy (ujjwal):

Prove that the max and min values of asinX + bcosX are respectively sqrt.(a^2+b^2) and -sqrt(a^2+b^2) .

OpenStudy (anonymous):

See -1<= sinx <=1 ==> -a<=asinx<=a -1<=cosx<=1 ==> -b<=bcosx <=b Now add them up -a-b<= asinx+bcosx <= a+b

OpenStudy (anonymous):

use compound angle formulae

OpenStudy (paxpolaris):

using derivatives: \[\Large a \cdot \cos(X) -b \cdot \sin(X)=0\]\[\huge \therefore \tan(X)=\frac ab\]

OpenStudy (experimentx):

\[ asinX + bcosX = \sqrt{a^2 + b^2} \cos (X - \tan^-1(b/a))\]

OpenStudy (experimentx):

put \( a = A\sin\phi \) and \( b = A\cos\phi \)

OpenStudy (anonymous):

i personally like @experimentX 's as it doesn't assume calculus

OpenStudy (anonymous):

but both are valid and nice :)

OpenStudy (experimentx):

lol ... not sure if this is called compound formula!!

OpenStudy (anonymous):

@eigenschmeigen ,@PaxPolaris ,@experimentX ,would you mind explaining why method won't work ??

OpenStudy (anonymous):

*my method

OpenStudy (paxpolaris):

 so, at max or min: \[y={a^2 \over \sqrt {a^2+b^2}} +{b^2 \over \sqrt {a^2+b^2}}\]

OpenStudy (experimentx):

well, @shivam_bhalla you assumed that x are different for sin and cos but they are same ... try for x=0 and x=pi/2

OpenStudy (paxpolaris):

** that should be \(\pm \sqrt{a^2+b^2}\) for the denominators

OpenStudy (anonymous):

Ok. getting it now @experimentX , Thanks a lot. I am still a noob in maths. :P

OpenStudy (ujjwal):

thanks everybody.. I got it...